Problem 10A-8 Applying Overhead; Overhead Variances [LO10-3, LO10-4] Lane Company manufactures a single product and applies oComplete this question by entering your answers in the tabs below. Req 2 Req 3A Req 1 Req 3B Req 4 Prepare a standard cost caComplete this question by entering your answers in the tabs below. Req 3B Req 1 Req 2 Req 3A Req 4 Compute the standard direcComplete this question by entering your answers in the tabs below. Req 3B Req 1 Req 2 Req 3A Req 4 Complete the following ManComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req 3A Req 3B Req 4 Determine the reason for a
Answer #1

Solution 1 and 2:

5 Solution 1: 6 | Computation of Predetermined overhead rate - Lane Company 7 Particulars Amount Budgeted variable manufactur

Solution 3a:

Standard hours allowed = 96000 units*1.5 hours = 144,000 hours

Solution 3b:

M Q solution 3b: Particulars Actual costs incurred(312000+936000) Manufacturing Overhead Debit | Particulars Credit $1,248,00

Solution 4:

Variable overhead actual rate = $312000/ 156000 = $2.00 per hour

Variable overhead rate variance = (SP - AP) *Actual Hours = ($3.20 - $2.00) * 156000 = $187,200 Favorable

Variable overhead Efficiency variance = (Standard hours - Actual Hours) *SP = (144000 - 156000)*$3.20 = $38,400 (Unfavorable)

Fixed overhead budget Variance = Budgeted Fixed Overhead - Actual Fixed overhead = $864000 - $936000 = $72,000 (unfavorable)

Fixed Overhead Volume Variance = ($7.20*144000) - $864000 = $172,800 Favorable